# Scalable unsupervised feature selection via weight stability

### v2

@@ 1 Introduction
Clustering is a fundamental machine learning technique that assigns data points in a dataset to clusters, with each cluster containing similar data points. As clustering algorithms operate without the need for labelled data, they have been successfully applied across various domains, including data pre-processing, quantitative finance, image analysis, and bioinformatics [1, 2, 3, 4]. However, their performance (and that of any other machine learning algorithm) can be heavily influenced by the quality and dimensionality of the data. High-dimensional datasets often contain redundant or irrelevant features, which can obscure meaningful patterns and degrade both efficiency and accuracy. Feature selection algorithms mitigate this issue by identifying and retaining only the most informative features [5, 6].
@@ 2.1 Clustering
The $k$-means algorithm [9] is a widely used partitional clustering method, which has been extended in numerous ways [13, 11]. It produces a partition $S=\{S_{1},\ldots,S_{k}\}$ of a dataset $X$ such that $S_{l}\cap S_{t}=\emptyset$ for all $l,t=1,\ldots k$ with $l\neq t$. This is achieved by minimising the objective function
@@
The $k$-means++ [14] algorithm addresses the latter of the problems above. It improves the initial centroids of $k$-means by spreading them out more strategically based on the data distribution. This often reduces the likelihood of poor local minima, leading to better clustering results. The $k$-means++ algorithm is now the default $k$-means implementation in many popular software packages such as MATLAB, scikit-learn, and R. Let $d(x_{i})$ be the distance from $x_{i}$ to its nearest centroid in $Z$. That is,
@@ 2.2 Unsupervised feature selection
Feature selection using feature similarity (FSFS) [20] is one of the most popular unsupervised feature selection methods (as indicated by its high number of citations in Google Scholar). Hence, we include it in our comparison (see Section 5). FSFS aims at identifying and removing redundant features by calculating their maximum information compression index,
@@
Multi-Cluster Feature Selection (MCFS) [21] is another popular unsupervised feature selection method. MCFS combines spectral techniques from manifold learning with $\ell_{1}$-regularization to identify the most relevant features. The algorithm has a single parameter $\kappa$, which specifies the number of neighbors in the $\kappa$-nearest neighbours graph. The original paper suggests setting $\kappa=5$ as a default value. In our experiments (see Section 5), we tuned this parameter from 1 to 5, and report the best results. While this procedure introduces a slight positive bias (i.e., it favours MCFS), it does not compromise our objectives. The algorithm is as follows.
@@
Subspace Clustering Feature Selection (SCFS) [22] is a recent unsupervised feature selection method combining subspace clustering with sparse regression. SCFS aims to select features that preserve the multi-cluster structure of the data by adaptively learning sample similarities. First, it computes a similarity matrix implicitly by solving a self-expressive model, in which each sample is represented through a low-dimensional space shared with other similar samples. Then, a regularised regression model is applied to link features to the discovered clustering structure. Features with stronger associations to the cluster are ranked higher. SCFS introduces two regularisation paramenters, $\alpha$ and $\beta$, to control the trade-off between similarity preservation and sparsity. Following the original paper we tuned these parameters using grid search over the set $\{10^{-4},10^{-2},1,10^{2},10^{4}\}$ and selected the best performing combination. We direct readers interested in further details about SCFS to its original publication [22].
Laplacian Score-regularized Concrete Autoencoder (LS-CAE) [23] is a recent deep learning-based approach to unsupervised feature selection. LS-CAE extends the Concrete Autoencoder (CAE) framework by incorporating a Laplacian score regularisation term into the loss function. This encourages the selection of features that both enable reconstruction of the original data and preserve the local clustering structure. To achieve this, LS-CAE trains a fully differentiable autoencoder, where feature selection is implemented through a Concrete layer that softly samples features in a learnable manner. The final selected features are obtained after annealing the sampling temperature towards discrete choices.

### v3

@@ 1 Introduction
Clustering is a fundamental machine learning technique that assigns data points in a dataset to clusters, with each cluster containing similar data points. As clustering algorithms operate without the need for labelled data, they have been successfully applied across various domains, including data pre-processing, quantitative finance, image analysis, and bioinformatics [1, 2, 3, 4, 5]. However, their performance (and that of any other machine learning algorithm) can be heavily influenced by the quality and dimensionality of the data. High-dimensional datasets often contain redundant or irrelevant features, which can obscure meaningful patterns and degrade both efficiency and accuracy. Feature selection algorithms mitigate this issue by identifying and retaining only the most informative features [6, 7].
@@ 2.1 Clustering
The $k$-means algorithm [10] is a widely used partitional clustering method, which has been used and extended in numerous ways [14, 15, 16, 12]. It produces a partition $S=\{S_{1},\ldots,S_{k}\}$ of a dataset $X$ such that $S_{l}\cap S_{t}=\emptyset$ for all $l,t=1,\ldots k$ with $l\neq t$. This is achieved by minimising the objective function
@@
The $k$-means++ [17] algorithm addresses the latter of the problems above. It improves the initial centroids of $k$-means by spreading them out more strategically based on the data distribution. This often reduces the likelihood of poor local minima, leading to better clustering results [18, 19]. The $k$-means++ algorithm is now the default $k$-means implementation in many popular software packages such as MATLAB, scikit-learn, and R. Let $d(x_{i})$ be the distance from $x_{i}$ to its nearest centroid in $Z$. That is,
@@ 2.2 Unsupervised feature selection
Feature selection using feature similarity (FSFS) [25] is one of the most popular unsupervised feature selection methods [26, 27, 28], which is also indicated by its high number of citations in Google Scholar. Hence, we include it in our comparison (see Section 5). FSFS aims at identifying and removing redundant features by calculating their maximum information compression index,
@@
Multi-Cluster Feature Selection (MCFS) [29] is another popular unsupervised feature selection method [6, 30]. MCFS combines spectral techniques from manifold learning with $\ell_{1}$-regularization to identify the most relevant features. The algorithm has a single parameter $\kappa$, which specifies the number of neighbors in the $\kappa$-nearest neighbours graph. The original paper suggests setting $\kappa=5$ as a default value. In our experiments (see Section 5), we tuned this parameter from 1 to 5, and report the best results. While this procedure introduces a slight positive bias (i.e., it favours MCFS), it does not compromise our objectives. The algorithm is as follows.
@@
Subspace Clustering Feature Selection (SCFS) [31] is a recent unsupervised feature selection method combining subspace clustering with sparse regression. SCFS aims to select features that preserve the multi-cluster structure of the data by adaptively learning sample similarities [32]. First, it computes a similarity matrix implicitly by solving a self-expressive model, in which each sample is represented through a low-dimensional space shared with other similar samples. Then, a regularised regression model is applied to link features to the discovered clustering structure. Features with stronger associations to the cluster are ranked higher. SCFS introduces two regularisation paramenters, $\alpha$ and $\beta$, to control the trade-off between similarity preservation and sparsity. Following the original paper we tuned these parameters using grid search over the set $\{10^{-4},10^{-2},1,10^{2},10^{4}\}$ and selected the best performing combination. We direct readers interested in further details about SCFS to its original publication [31].
Laplacian Score-regularized Concrete Autoencoder (LS-CAE) [33] is a recent deep learning-based approach to unsupervised feature selection [26]. LS-CAE extends the Concrete Autoencoder (CAE) framework by incorporating a Laplacian score regularisation term into the loss function. This encourages the selection of features that both enable reconstruction of the original data and preserve the local clustering structure. To achieve this, LS-CAE trains a fully differentiable autoencoder, where feature selection is implemented through a Concrete layer that softly samples features in a learnable manner. The final selected features are obtained after annealing the sampling temperature towards discrete choices.